# A Systematic Literature Review of Software Engineering Research on Jupyter Notebook

### v1

@@ 3.1.1 Identifying the Need for a Systematic Review
The need for a systematic review of software engineering research on Jupyter Notebook arises from the growing popularity and widespread adoption of Jupyter notebooks in various research fields. The widespread use of Jupyter notebooks has highlighted software engineering challenges such as poor code quality, lack of coding standards, and code duplication [1, 179, 73]. These difficulties have sparked notable interest among software engineering researchers in helping notebook users adopt software engineering practices in their work. As a result, more research has been conducted on these issues in Jupyter notebooks. This highlights the need to conduct a systematic literature review of this research and the challenges it addresses, as well as identify what should be studied next.
Figure 2: The steps of our SLR on software engineering research on Jupyter Notebook based on Kitchenham’s guidelines [71]
@@ 3.1.2 Developing the Review Protocol
A review protocol is necessary to outline the procedure for conducting the systematic review to limit the likelihood of researcher bias [71]. To meet these objectives, our protocol involves: (1) defining research questions, (2) searching and selecting primary studies, (3) extracting data from primary studies, and (4) synthesizing the data.
@@
    Motivation: The first research question gives a quantitative overview of the software engineering research on Jupyter Notebook. RQ1 aims to provide a high-level overview of how the field of software engineering research on Jupyter Notebook is evolving and where it is getting attention. RQ1 also explores what kinds of solutions are being proposed by software engineering researchers to deal with problems in Jupyter notebooks. By answering RQ1, we can identify trends and patterns of software engineering research targeted at Jupyter notebooks over time. Furthermore, this question explores whether research in this area follows best practices for open science [98].
@@ 3.2.1 Searching and Selecting Primary Studies
We searched all the indexed articles in the DBLP  [^8] https://dblp.org/ to find relevant papers for this systematic review. DBLP is a computer science bibliography website that provides free access to bibliographic information on major computer science publications. It includes conference papers, journal articles, and other academic reports related to computer science. It is a reputable and comprehensive source of academic papers, and many researchers [193, 32, 82] used it to find relevant studies for their systematic literature reviews.
The focus of our review is software engineering research on Jupyter Notebook. This type of research includes work that analyzes or improves current software engineering practices in notebooks. To find relevant studies for our review, we searched the title of the papers with the following query: “Notebook” or “Jupyter”. The literature list was compiled up to the end of 2024, capturing the latest software engineering research on Jupyter notebooks.
@@ 3.2.3 Synthesizing the Data
We manually explored the title of the online version of all those papers. In this stage, we excluded several studies that did not focus on computational notebooks but on physical notebooks (e.g., laptops). Then, we explored the studies’ abstracts and selected the papers that target software engineering-related topics on Jupyter Notebook. We also excluded studies that focused on topics that are adjacent to software engineering research on Jupyter Notebook, such as those on using Jupyter Notebook for software engineering education. For example, we excluded the experience report by Al-Gahmi et al. [4] on using Jupyter Notebook in classroom programming.
We manually reviewed the list of studies to exclude duplicates, e.g., sometimes a study’s final official version and unofficial pre-print version are available online. For example, Wang et al. published a pre-print of their paper titled “Themisto: Towards Automated Documentation Generation in Computational Notebooks” [173]. After that, the same paper was published officially in the next year with a different title “Documentation Matters: Human-Centered AI System to Assist Data Science Code Documentation in Computational Notebooks” [174]. We manually checked both versions of that paper and excluded the earlier version.
To ensure the reliability of our study selection procedure, the first and the third authors independently went through the list of studies following the procedure above. We used Cohen’s Kappa statistic to measure the Inter-Rater Reliability (IRR) [96], which indicates the level of agreement between two raters in a classification task. We found that Cohen’s Kappa value is 0.79, which indicates substantial agreement between the raters. We measured the IRR into two stages. In the first stage, the first and third authors discussed and resolved each disagreement to come up with a list of selected studies up to 2023. In the second stage, studies from 2024 were discussed and resolved by the first and second authors. The disagreements mostly arose from the first author’s misunderstanding of defining the scope of provenance in the notebook. After discussing and resolving these disagreements, the Kappa value became 1.0 after resolving the disagreements. As a result, we ended up with a list of 146 studies for our SLR.
@@ 4.1 Year-wise Distributions
The data from the SLR showcases a marked increase in software engineering studies on Jupyter notebooks. The graph in Figure 3 shows the number of publications with software engineering research on Jupyter notebooks per year. The figure outlines a growing trend. Software engineering research on notebooks began with just one study in 2015 and followed an increasing trend to 42 in 2024. This trend reflects a growth in academic interest in considering software engineering practices in Jupyter notebooks.
@@ 4.2 Publication Types
The majority of the work on software engineering research on notebooks was published at conferences. Figure 4 shows the number of studies per type. Our research findings showed that the majority, specifically 53.4% (78 out of 146 studies), were presented at conferences. In addition, 26 studies were published in journals, 11 in workshops, and 11 in symposia. Furthermore, we identified 20 studies that appeared in informal and other types of publications, such as book chapters, thesis dissertations, or those published in the arXiv repository. It is important to note that conferences publish findings more quickly than journals, allowing researchers to share updates with the community faster [44]. The substantial number of conference papers included in the systematic literature review indicates that software engineering research on notebooks is an active research area of study.
@@ 4.3 Publication Venues
The most popular venues for software engineering research on notebooks are focused on human-computer interaction (HCI). Table LABEL:tab:journalSymnWorkshopOthers presents the list of conferences, journals, symposiums, workshops, and other publication venues where the primary studies have been published. We found that the most popular venue (15 studies) was the ACM Conference on Human Factors in Computing Systems (CHI). Likewise, the most popular symposium (six studies) is the IEEE Symposium on Visual Languages and Human-Centric Computing (VL/HCC), and one of the most popular journals (two studies) is the ACM Transactions on Computer-Human Interaction (TOCHI). This pattern emphasizes the interdisciplinary nature of working with Jupyter notebooks and underlines the importance of their user experience and usability (which are important topics at the HCI venues above).
@@
Conference CHI ACM Conference on Human Factors in Computing Systems [55, 64, 84, 81, 97, 125, 135, 172, 183, 189, 21, 54, 60, 175, 184] 15
ASE Automated Software Engineering Conference [10, 30, 100, 177, 197, 203, 59] 7
ICSE International Conference on Software Engineering [109, 119, 155, 180, 178, 179] 6
MSR Mining Software Repositories Conference [46, 120, 115, 40] 4
@@
TaPP USENIX Conference on Theory and Practice of Provenance. [28, 114, 75, 117] 4
APSEC Asia-Pacific Software Engineering Conference [131, 145, 154] 3
SANER Int. Conference on Software Analysis, Evolution and Reengineering [159, 167] 2
CAIN Int. Conference on AI Engineering – Software Engineering for AI [122, 41] 2
EDBT International Conference on Extending Database Technology [56, 186] 2
@@
Conference e-Science IEEE International Conference on eScience [3, 25] 2
CIDR Int. Conference on Innovative Data Systems Research [93, 12] 2
SciPy Scientific Computing with Python Conference [99, 92] 2
ICSME Int. Conference on Software Maintenance and Evolution [61, 205] 2
FSE ACM International Conference on the Foundations of Software Engineering [29, 181] 2
IUI Annual Conference on Intelligent User Interfaces [101, 38] 2
SCAM IEEE Working Conference on Source Code Analysis and Manipulation [149] 1
EASE Int. Conference on Evaluation and Assessment in Software Engineering [2] 1
@@
PEARC Practice and Experience in Advanced Research Computing Conference [185] 1
UCC International Conference on Utility and Cloud Computing [196] 1
EMNLP Int. Conference Empirical Methods in Natural Language Processing [91] 1
SPLC Int. Systems and Software Product Line Conference [13] 1
IJCAI International Joint Conference on Artificial Intelligence [169] 1
FLAIRS Int. Florida Artificial Intelligence Research Conference [105] 1
WWW ACM Web Conference [77] 1
ACL Annual Meeting of the Association for Computational Linguistics [201] 1
@@
ICPC IEEE/ACM International Conference on Program Comprehension [132] 1
AIMLS International Conference on AI ML Systems [190] 1
Programming ACM Conference on the Art, Science, and Engineering of Programming [104] 1
KDIR Int. Conference on Knowledge Discovery, Engineering, and Management [69] 1
ISWC International Semantic Web Conference [137] 1
@@
ICMD International Conference on Management of Data [130] 1
QCE International Conference on Quantum Computing and Engineering [70] 1
SEAA Euromicro Conference on Software Engineering and Advanced Applications [43] 1
CSR International Conference on Cyber Security and Resilience [128] 1
MOD International Conference on Management of Data [86] 1
JCSSE Int. Joint Conference on Computer Science and Software Engineering [126] 1
Journal EMSE Empirical Software Engineering Journal [116, 127, 166] 3
TOCHI ACM Transactions on Computer-Human Interaction [121, 174] 2
ACM-HCI Proceedings of the ACM on Human-Computer Interaction [134, 187] 2
VLDB International Journal on Very Large Data Bases [88, 146] 2
SoftwareX SoftwareX [35, 123] 2
IEEE TVCG IEEE Transactions on Visualization and Computer Graphics [89, 144] 2
TOSEM ACM Transactions on Software Engineering and Methodology [90, 26] 2
ACM PL Proceedings of the ACM on Programming Languages [141] 1
Sigmod Rec. ACM SIGMOD Record Journal [118] 1
CGF Computer Graphics Forum [37] 1
Interactions ACM Interactions [124] 1
GigaScience GigaScience [139] 1
BTW Journal on Business, Technologie und Web [68] 1
JASEP The Journal on the Art, Science, and Engineering of Programming [63] 1
PLoS CB. PLOS Computational Biology Journal [133] 1
TiiS ACM Transactions on Interactive Intelligent Systems [85] 1
SoftwareTT Softwaretechnik-Trends Journal [152] 1
JASIST Journal of the Association for Information Science and Technology [15] 1
@@
Symposium VL/HCC Symposium on Visual Languages and Human-Centric Computing [73, 22, 156, 78, 66, 23] 6
@@
UIST ACM Symposium on User Interface Software and Technology [195, 107] 2
@@
AISTA ACM International Symposium on Software Testing and Analysis [165] 1
@@
Onward ACM SIGPLAN International Symposium on New Ideas, New Paradigms, and Reflections on Programming and Software [151] 1
SPLASH-E ACM SIGPLAN International Symposium on SPLASH-E [76] 1
@@
Workshop IDE ACM/IEEE Workshop on Integrated Development Environments [170, 160] 2
@@
IPAW International Provenance and Annotation Workshop [138, 74] 2
@@
SC Workshop Workshops of the International Conference for High Performance Computing, Networking, Storage and Analysis [16, 191] 2
SOAP ACM Int. Workshop on the State Of the Art in Program Analysis [103] 1
@@
QuASoQ International Workshop on Quantitative Approaches to Software Quality [7] 1
@@
HILDA Workshop on Human-In-the-Loop Data Analytics [14] 1
@@
PAINT Workshop on Programming Abstractions and Interactive Notations, Tools, and Environments [168] 1
IWSC International Workshop on Software Clones [140] 1
Informal and Other Publications arXiv An archive for electronic preprints of scientific studies [20, 24, 31, 34, 39, 45, 47, 50, 57, 87, 83, 111, 143, 147, 182, 188, 194, 198, 204] 19
@@
Thesis Thesis dissertation was published at the University of California, San Diego [151] 1

### v2

@@ 3.1.1 Identifying the Need for a Systematic Review
The need for a systematic review of software engineering research on Jupyter Notebook arises from the growing popularity and widespread adoption of Jupyter notebooks in various research fields. The widespread use of Jupyter notebooks has highlighted software engineering challenges such as poor code quality, lack of coding standards, and code duplication [adams2023comparison, wang2020better, koenzen2020code]. These difficulties have sparked notable interest among software engineering researchers in helping notebook users adopt software engineering practices in their work. As a result, more research has been conducted on these issues in Jupyter notebooks. This highlights the need to conduct a systematic literature review of this research and the challenges it addresses, as well as identify what should be studied next.
Figure 2: The steps of our SLR on software engineering research on Jupyter Notebook based on Kitchenham’s guidelines [kitchenham2004procedures]
@@ 3.1.2 Developing the Review Protocol
A review protocol is necessary to outline the procedure for conducting the systematic review to limit the likelihood of researcher bias [kitchenham2004procedures]. To meet these objectives, our protocol involves: (1) defining research questions, (2) searching and selecting primary studies, (3) extracting data from primary studies, and (4) synthesizing the data.
@@
    Motivation: The first research question gives a quantitative overview of the software engineering research on Jupyter Notebook. RQ1 aims to provide a high-level overview of how the field of software engineering research on Jupyter Notebook is evolving and where it is getting attention. RQ1 also explores what kinds of solutions are being proposed by software engineering researchers to deal with problems in Jupyter notebooks. By answering RQ1, we can identify trends and patterns of software engineering research targeted at Jupyter notebooks over time. Furthermore, this question explores whether research in this area follows best practices for open science [Mendez_2020].
@@ 3.2.1 Searching and Selecting Primary Studies
We searched all the indexed articles in IEEE Xplore, ACM Digital Library, Elsevier ScienceDirect, Springer Nature Link, and DBLP. Together, these five sources comprehensively cover conference papers, journal articles, and other academic reports related to computer science and software engineering. The focus of our review is software engineering research on Jupyter Notebook. This type of research includes work that analyzes or improves current software engineering practices in notebooks. To find relevant studies for our review, we searched the title of the papers with the following query: “Notebook” or “Jupyter”. The literature list was compiled up to September 2025, capturing the latest software engineering research on Jupyter notebooks.
@@ 3.2.3 Synthesizing the Data
We manually explored the title of the online version of all those papers. In this stage, we excluded several studies that did not focus on computational notebooks but on physical notebooks (e.g., laptops). Then, we explored the studies’ abstracts and selected the papers that target software engineering-related topics on Jupyter Notebook. We also excluded studies that focused on topics that are adjacent to software engineering research on Jupyter Notebook, such as those on using Jupyter Notebook for software engineering education. For example, we excluded the experience report by al2022jupyter on using Jupyter Notebook in classroom programming.
We manually reviewed the list of studies to exclude duplicates, e.g., sometimes a study’s final official version and unofficial pre-print version are available online. For example, Wang et al. published a pre-print of their paper titled “Themisto: Towards Automated Documentation Generation in Computational Notebooks” [wang2021themisto]. After that, the same paper was published officially in the next year with a different title “Documentation Matters: Human-Centered AI System to Assist Data Science Code Documentation in Computational Notebooks” [wang2022documentation]. We manually checked both versions of that paper and excluded the earlier version.
To ensure the reliability of our study selection procedure, the first and third authors independently went through the list of studies following the procedure above. We used Cohen’s Kappa statistic to measure the Inter-Rater Reliability (IRR) [mchugh2012interrater], which indicates the level of agreement between two raters in a classification task. We measured the IRR into two stages. In the first stage, the first and third authors discussed and resolved each disagreement to come up with a list of selected studies up to 2023. In the second stage, studies after 2023 and till September 2025 were discussed and resolved by the first and second authors. The disagreements mostly arose from the first author’s misunderstanding of defining the scope of provenance in the notebook. The Cohen’s Kappa value was 0.73 for the first stage and 0.97 for the second stage, indicating substantial agreement in both stages. After resolving all disagreements, we identified 195 papers. We then conducted backward and forward snowballing, identifying four additional studies. Two of these had been unintentionally missed during the labeling process, which demonstrates the overall robustness of our selection procedure, as only two studies were overlooked. The other two were newly published after our cutoff date and were not yet indexed in the source databases. In total, we finalized 199 studies as the primary set for this SLR.
@@ 4.1 Year-wise Distributions
The data from the SLR showcases a marked increase in software engineering studies on Jupyter notebooks. The graph in Figure 3 shows the number of publications with software engineering research on Jupyter notebooks per year. The figure outlines a growing trend. Software engineering research on notebooks began with just one study in 2015 and followed an increasing trend to 47 in 2024 and 31 in 2025 (until September 25th). This trend reflects a growth in academic interest in considering software engineering practices in Jupyter notebooks.
@@ 4.2 Publication Types
The majority of the work on software engineering research on notebooks was published at conferences. Figure 4 shows the number of studies per type. Our research findings showed that the majority, specifically 52.3% (104 out of 199 studies), were presented at conferences. In addition, 37 studies were published in journals, 17 in symposia, and 14 in workshops. Furthermore, we identified 27 studies that appeared in informal and other types of publications, such as book chapters, thesis dissertations, or those published in the arXiv repository. It is important to note that conferences publish findings more quickly than journals, allowing researchers to share updates with the community faster gonzalez2011articles. The substantial number of conference papers included in the systematic literature review indicates that software engineering research on notebooks is an active research area of study.
@@ 4.3 Publication Venues
The most popular venues for software engineering research on notebooks are focused on human-computer interaction (HCI). Table LABEL:tab:journalSymnWorkshopOthers presents the list of conferences, journals, symposiums, workshops, and other publication venues where the primary studies have been published. We found that the most popular venue (20 studies) was the ACM Conference on Human Factors in Computing Systems (CHI). Likewise, the most popular symposium (eight studies) is the IEEE Symposium on Visual Languages and Human-Centric Computing (VL/HCC), and one of the popular journals (three studies) is the Springer International Journal on Very Large Data Bases (VLDB Journal). This pattern emphasizes the interdisciplinary nature of working with Jupyter notebooks and underlines the importance of their user experience and usability (which are important topics at the HCI venues above).
@@
Conference CHI ACM Conference on Human Factors in Computing Systems head2019managing, kang2021toonnote, li2021nbsearch, li2023notable, mcnutt2023design, raghunandan2022code, rule2018exploration, wang2021makes, wang2022stickyland, weinman2021fork, chattopadhyay2020whats, harrison2024jupyterlab, in2024evaluating, wang2024outlinespark, wang2023supernova, fang2025enhancing, fang2025large, lin2025interlink, rawn2025pagebreaks, kery2018story 20
ASE Automated Software Engineering Conference bavishi2021vizsmith, dong2021splitting, mondal2023cell2doc, wang2020assessing, yang2022data, zhu2021restoring, huang2024code, mostafavi2024can 8
ICSE International Conference on Software Engineering patra2022nalin, quaranta2022assessing, subotic2022static, wang2021restoring, wang2020restoring, wang2020better, huang2025scientists 7
MSR Mining Software Repositories Conference grotov2022large, quaranta2021kgtorrent, pimentel2019large, mostafavi2024distilkaggle, nakamaru2025jupyter, nguyen2025majority 6
@@
TaPP USENIX Conference on Theory and Practice of Provenance. dao2022runtime, petricek2018wrattler, koop2017dataflow, pimentel2015collecting 4
@@
e-Science IEEE International Conference on eScience ahmad2022reproducible, cunha2021context, li2022context, li2023dense 4
@@
PEARC Practice and Experience in Advanced Research Computing Conference wannipurage2022framework, jayawardana2024enhancing, kim2024japper 3
@@
ICPC IEEE/ACM International Conference on Program Comprehension robinson2022error, husak2023slicito, wong2025method 3
APSEC Asia-Pacific Software Engineering Conference ritta2022reusing, settewong2022visualize, studtmann2023histree 3
FSE ACM International Conference on the Foundations of Software Engineering dong2021qualitative, wang2024using, jin2025learning 3
SciPy Scientific Computing with Python Conference michael2022keeping, lu2020securing 2
SANER Int. Conference on Software Analysis, Evolution and Reengineering titov2022resplit, venkatesh2023enhancing 2
IUI Annual Conference on Intelligent User Interfaces muller2021data, george2024notebookgpt 2
ICSME Int. Conference on Software Maintenance and Evolution jiang2022elevating, zou2024mlpipeline 2
EDBT International Conference on Extending Database Technology horiuchi2022jupysim, watson2019pysnippet 2
CIDR Int. Conference on Innovative Data Systems Research macke2021automating, brachmann2020your 2
CAIN Int. Conference on AI Engineering – Software Engineering for AI quaranta2022pynblint, mostafavi2024beyond 2
@@
ICMD International Conference on Management of Data li2024demonstration, rehman2019towards 2
WWW ACM Web Conference kwon2023weedle 1
@@
VL/HCC IEEE Symposium on Visual Languages and Human-Centric Computing harden2022exploring 1
VDS IEEE Visualization in Data Science (VDS) wenskovitch2019albireo 1
UCC International Conference on Utility and Cloud Computing xin2022provenance 1
SPLC Int. Systems and Software Product Line Conference brault2023taming 1
Conference SPLASH ACM SIGPLAN conference on Systems, Programming, Languages, and Applications: Software for Humanity christophe2023moon 1
ASSET International ACM SIGACCESS Conference on Computers and Accessibility christophe2023moon 1
SEAA Euromicro Conference on Software Engineering and Advanced Applications golendukhina2024unveiling 1
SCAM IEEE Working Conference on Source Code Analysis and Manipulation siddik2023code 1
QCE International Conference on Quantum Computing and Engineering kinanen2024improving 1
Programming ACM Conference on the Art, Science, and Engineering of Programming niephaus2019polyjus 1
PASC ACM Platform for Advanced Scientific Computing Conference tsai2024libyt 1
KDIR Int. Conference on Knowledge Discovery, Engineering, and Management kerzel2021towards 1
JCSSE Int. Joint Conference on Computer Science and Software Engineering ragkhitwetsagul2024typhon 1
ISWC International Semantic Web Conference samuel2018provbook 1
IJCAI International Joint Conference on Artificial Intelligence wang2021graph 1
ICIIBMS IEEE International Conference on Intelligent Informatics and Biomedical Sciences rehman2019towards 1
HCI International Conference on Human-Computer Interaction [schmieg2025enhancing] 1
FLAIRS Int. Florida Artificial Intelligence Research Conference oli2021automated 1
EMNLP Int. Conference Empirical Methods in Natural Language Processing liu2021haconvgnn 1
EDUCON IEEE Education Engineering [bajaj2025tool] 1
EASE Int. Conference on Evaluation and Assessment in Software Engineering agrawal2022understanding 1
CSR International Conference on Cyber Security and Resilience ramsingh2024understanding 1
CSCL International Conference on Computer-Supported Collaborative Learning cai2025jupyterNotebook 1
Coil Calling ACM Symposium on Eye Tracking Research and Applications [sparmann2023jugaze] 1
AIMLS International Conference on AI ML Systems werner2021bridging 1
ACL Annual Meeting of the Association for Computational Linguistics yin2023natural 1
Journal EMSE Empirical Software Engineering Journal pimentel2021understanding, ramasamy2023visualising, venkatesh2024static, ghahfarokhi2024predicting 4
VLDB International Journal on Very Large Data Bases li2023elasticnotebook, shankar2022bolt, zhang2019juneau 3
TOSEM ACM Transactions on Software Engineering and Methodology liu2023refactoring, de2024bug, wang2024why 3
Journal IEEE TVCG IEEE Transactions on Visualization and Computer Graphics lin2023inksight, scully2024design, wootton2024charting 3
TOCHI ACM Transactions on Computer-Human Interaction quaranta2022eliciting, wang2022documentation 2
ACM-HCI Proceedings of the ACM on Human-Computer Interaction rule2018aiding, weber2024extending 2
SoftwareX SoftwareX fitzpatrick2022davos, quaranta2024pynblint 2
CGF Wiley Computer Graphics Forum gadhave2024persist, nylund2025matplotalt 2
ACM PL Proceedings of the ACM on Programming Languages sato2024multiverse 1
Sigmod Rec. ACM SIGMOD Record Journal psallidas2022data 1
Interactions ACM Interactions raghunandan2023measuring 1
GigaScience GigaScience samuel2022computational 1
BTW Journal on Business, Technologie und Web kerzel2023mlprovlab 1
JASEP The Journal on the Art, Science, and Engineering of Programming kallen2021jupyter 1
PLoS CB. PLOS Computational Biology Journal rule2018ten 1
TiiS ACM Transactions on Interactive Intelligent Systems li2023edassistant 1
SoftwareTT Softwaretechnik-Trends Journal speicher2019notes 1
JASIST Journal of the Association for Information Science and Technology candela2023approach 1
Softw. Pract. Exp. Wiley Software: Practice and Experience li2025search 1
GIB Software Engineering: Gesellschaft für Informatik, Bonn mietchen2025analyzing 1
JGI Journal of Graphics Interface harden2023there 1
JCSE Journal of Computing in Science and Engineering ono2021interactive 1
TVCG IEEE Trans. Visualization and Computer Graphics eckelt2024loops 1
CCSC Journal of Computing Sciences in Colleges adams2023comparison 1
Symposium VL/HCC Symposium on Visual Languages and Human-Centric Computing koenzen2020code, chen2023whatsnext, subramanian2020casual, lau2020design, kery2018interactions, cheng2024biscuit, christophe2023moon, merino2022making 8
@@
UIST ACM Symposium on User Interface Software and Technology wu2020b2, ouyang2024noteplayer 2
@@
AISTA ACM International Symposium on Software Testing and Analysis venkatesh2021automated 1
@@
Onward ACM International Symposium on New Ideas, New Paradigms, and Reflections on Programming and Software singer2020notes 1
SPLASH-E ACM SIGPLAN International Symposium on SPLASH-E kimio2024kogi 1
SIGCSE ACM Technical Symposium on Computer Science Education cai2025jupyter 1
NOMS IEEE/IFIP Network Operations and Management Symposium ueno2020ssh 1
NCA International Symposium on Network Computing and Applications faenza2024containerized 1
LDAV IEEE Symposium on Large Data Analysis and Visualization savira2021writing 1
@@
Workshop IDE ACM/IEEE Workshop on Integrated Development Environments titov2024hidden, wang2024dont 2
@@
IPAW International Provenance and Annotation Workshop koop2021notebook, samuel2021reproducemegit 2
@@
SC Workshop Workshops of the International Conference for High Performance Computing, Networking, Storage and Analysis cao2024jupyter, werner2024jumper 2
SOAP ACM Int. Workshop on the State Of the Art in Program Analysis negrini2023static 1
@@
QuASoQ International Workshop on Quantitative Approaches to Software Quality aydin2024cellrecommend 1
@@
HILDA Workshop on Human-In-the-Loop Data Analytics brown2023facilitating 1
@@
PAINT Workshop on Programming Abstractions and Interactive Notations, Tools, and Environments verano2022suppose 1
IWSC International Workshop on Software Clones sato2022comparing 1
IPDPSW International Parallel and Distributed Processing Symposium Workshops oden2024integrating 1
DL4Code International Workshop on Deep Learning for Code grotov2025themisto 1
IDE IEEE/ACM Second IDE Workshop grotov2025evolving 1
Informal and Other Publications arXiv An archive for electronic preprints of scientific studies carver2025notebookos, chattopadhyay2023make, choetkiertikul2023mining, duan2023jup2kub, fangohr2020testing, grotov2024debug, grotov2024untangling, guo2024explainability, horiuchi2022similarity, in2025exploring, jiang2025exploring, jin2025suggesting, li2024kishu, li2024unlocking, li2025kernel, perez2024flexible, schroder2019reproducible, shome2024understanding, tang2025characterising, tian2025noteflow, titov2025observing, wandel2025pyevalai, weber2024computational, yang2023code, you2025datawiseagent, zhu2024facilitating 26
@@
Thesis Thesis dissertation was published at the University of California, San Diego singer2020notes 1